OpenStudy (anonymous):

What is the measure of angle C? (Points : 2) 28° 62° 90° 10°

OpenStudy (anonymous):

@jobs

OpenStudy (nurali):

A..not for sure on that one.

OpenStudy (anonymous):

@jayz657

OpenStudy (anonymous):

6x+30 + 5x + 12 + 2x + 8 = 180 13x + 50 = 180 13x = 130 x = 10 so angle c is 2(10) + 8 = 28 degrees

OpenStudy (anonymous):

6x+30 = 90 6x = 60 x= 10 2(10)+ 8 = 28 degrees ----------->required